OpenStudy (anonymous):

If y2 = x3 + 2, which statement is true? A. The equation is a function at all values of x. B. The equation is not a function. C. The equation is a function only when x < 0 D. The equation is a function only when x > 0.

OpenStudy (anonymous):

A) is a function at all values of x.

OpenStudy (michele_laino):

hint: our function exixts if and only if: \[x + 2 \geqslant 0\]

OpenStudy (michele_laino):

exists*

OpenStudy (michele_laino):

more precisely the expression: \[\sqrt {{x^3} + 2} \] exists if: \[x + 2 \geqslant 0\]

OpenStudy (michele_laino):

we have to keep in mind that we have two branches, namely: \[y = \sqrt {{x^3} + 2} \] and \[y = - \sqrt {{x^3} + 2} \]

OpenStudy (michele_laino):

oops..I have made an error, since the two branches esist if \[x + \sqrt[3]{2} \geqslant 0\]

OpenStudy (michele_laino):

we have this graph, approximatively  now please make the vertical line test